OpenStudy (anonymous):

Given segment AB with endpoints A(2x, 4y) and B(10, 22), what is the midpoint? A.(2x+5, 4y+11) B.(x+2 ½ , y+5 ½) C.(2x-10, 4y-11) D.(x+5, 2y+11) I know the midpoint formula, but it doesnt seem to work out right. Can someone explain it to me...

OpenStudy (anonymous):

\[\frac{ x1 + x2 }{ 2 }, \frac{ y1 + y2 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 10 + 2x }{ 2 }, \frac{ 22 + 4y }{ 2 }\] now simplify those fractions

OpenStudy (anonymous):

ok thanks. the only thing that confuses me is the variables. but let me try it real fast ...

OpenStudy (anonymous):

you can look at it like \[\frac{ 10 + 2 }{ 2 }\] then just remember to add the x back in on the top

OpenStudy (anonymous):